OpenStudy (anonymous):

how do i solve a problem like this 2x-25

OpenStudy (anonymous):

2x-25 = (x+5)(x-5) x=-5, +5

OpenStudy (anonymous):

Thank you!!!! this helped out a lot

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

@redred413 can you explain how she did this please!

OpenStudy (anonymous):

nada bit sorry :( im not good at math

OpenStudy (anonymous):

ahh alright lol

OpenStudy (anonymous):

@.Sam. hey can you show me how she did this problem?

OpenStudy (anonymous):

@phi hey can you show me how she did this problem?

OpenStudy (anonymous):

@nubeer @chmvijay hey can you one of you show me how she solved the problem?

OpenStudy (nubeer):

\[2x-25 \] is this your question?

OpenStudy (anonymous):

yes!

OpenStudy (nubeer):

or \[x ^{2}-25\]

OpenStudy (anonymous):

no the first one was right 2x-25

OpenStudy (nubeer):

2x -25 =0 1) add 25 on both sides. 2) divide both sides by 2

OpenStudy (chmvijay):

LOL she mentioned two factors for X not one if you do by ur method only +5 will get what about -5

OpenStudy (phi):

***how do i solve a problem like this 2x-25**** this is not solvable unless it is an equation if you had 2x-25=0 you can find the x that makes it true. see post by nubeer for details

OpenStudy (phi):

if the expression was \[ x^2 -25 \] you could not "solve it", but you could factor it... write it in a different way that means the same thing. This would be a "difference of squares" and you would use the rule \[ (a^2 - b^2)= (a-b)(a+b) \] for this problem you would do this \[ x^2 -25 = (x-5)(x+5) \]

OpenStudy (anonymous):

ohh so it is pretty much i just need find the GCF

OpenStudy (anonymous):

and use the equation